OpenStudy (alina123):

A recipe calls for trail mix for 3/4 cup of dried fruit, 1/2 cup of mixed nuts, and 1/3 cup of granola. How many cups of trail mix does this recipe make? * Draw a model o answer this question *Write and evaluate a numerical expression to answer the question. * Explain how you know your answer is correct?

OpenStudy (e.mccormick):

Get all the fractions to a common denominator.

OpenStudy (alina123):

so you add them up together than

OpenStudy (e.mccormick):

Yep.

OpenStudy (alina123):

yeh but what does that give the whole trail mix

OpenStudy (alina123):

and the how many cups

OpenStudy (e.mccormick):

Yes, it would give the total as a fraction in cups. You just need to simplify the fraction, if possible, to get the best form of the answer.

OpenStudy (texaschic101):

yes. You will arrive at an improper fraction which you will turn to a mixed number.

OpenStudy (alina123):

oh ok and so what is my expresion

OpenStudy (texaschic101):

3/4 + 1/2 + 1/3 = total cups

OpenStudy (alina123):

so thats my numerical expression

OpenStudy (texaschic101):

well.....thats more of an equation....I guess your expression would be 3/4 + 1/2 + 1/3

OpenStudy (anonymous):

Like so; 3/4 + 1/2 + 1/3

OpenStudy (alina123):

ok and so my diagram is what

OpenStudy (alina123):

diagram or model

OpenStudy (anonymous):

would be 9/12 + 6/12 + 4/12 = 19/12

OpenStudy (alina123):

that my model

OpenStudy (texaschic101):

it probably has to be drawn

OpenStudy (alina123):

of what the mixes

OpenStudy (anonymous):

simplify, 19 divided by 12 (about equals)= 1.58

OpenStudy (texaschic101):

is drawn a word ? You probably have to draw it...thats better

OpenStudy (anonymous):

about 1 and a half cups! voila! :)

OpenStudy (alina123):

um potami y do we need decimals

OpenStudy (texaschic101):

19/12 = 1 7/12 cups

OpenStudy (anonymous):

@Alina123 it makes it easier =)

OpenStudy (alina123):

sorry for the name

OpenStudy (alina123):

and ok

OpenStudy (alina123):

oh i get it, k thxs everybody!

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

np! happy to help, and that's ok (for the name)

OpenStudy (alina123):

Guys my teacher said I got it wrong why

OpenStudy (e.mccormick):

If they did not explain the mistake it is impossible to know why it was seen as wrong. For example, you may have had the right numbers but they may have felt the model was done in the wrong way.